botany500:

The graph shows the height, y, in centimeters, of a plant after a certain number of weeks, x. Linda drew the line of best fit on the graph. A graph titled Plant Height shows Number of Weeks on x axis and Height of Plant in cm on y axis. The scales on both x and y axes are shown from 0 to 5 at increments of 5. The graph shows dots at the ordered pairs 0, 2 and 1, 2.4 and 1.5, 2.7 and 2, 2.8 and 2.5, 3 and 3, 3.2 and 3.5, 3.5 and 4, 3.5 and 4.5, 3.6 and 5, 4. A straight line joins the ordered pairs 0, 2 and 5, 4 What would most likely be the approximate height of the plant after 7 weeks?

botany500:

please help hero?

Hero:

Hint: For points (0,2) and (5,4) Use the two point formula to find the equation of a line: \(y - y_1 = \dfrac{y_2 - y_1}{y_2 - y_1}(x - x_1)\)

botany500:

I do not get it

misty34:

Y2 is 2 and y1 0?

misty34:

I forgot the order they go but, something like that

Hero:

No, that's not right @misty34

Hero:

\((x_1,y_1) = (0,2)\) \((x_2,y_2) = (5,4)\)

misty34:

I was on the right track but, forgot the order

misty34:

It's been a while since I've done those

misty34:

like 4-2 is 1 of them?

misty34:

then 5-0

Hero:

Actually the formula is \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}(x - x_1)\)

botany500:

is it 4.8

Hero:

And the result will be a linear equation of the form \(y = mx + b\)

misty34:

I have the numbers mixed up again

botany500:

y=2m+5

Hero:

In the resulting equation, m has a value.

Hero:

x and y will be the variables